# When Tokens Talk Too Much: A Survey of Multimodal Long-Context Token Compression across Images, Videos, and Audios

### v3

@@ Abstract
Multimodal large language models (MLLMs) have made remarkable strides, largely driven by their ability to process increasingly long and complex contexts, such as high-resolution images, extended video sequences, and lengthy audio input. While this ability significantly enhances MLLM capabilities, it introduces substantial computational challenges, primarily due to the quadratic complexity of self-attention mechanisms with numerous input tokens. To mitigate these bottlenecks, token compression has emerged as an auspicious and critical approach, efficiently reducing the number of tokens during both training and inference. In this paper, we present the first systematic survey and synthesis of the burgeoning field of multimodal long context token compression. Recognizing that effective compression strategies are deeply tied to the unique characteristics and redundancies of each modality, we categorize existing approaches by their primary data focus, enabling researchers to quickly access and learn methods tailored to their specific area of interest: (1) image-centric compression, which addresses spatial redundancy in visual data; (2) video-centric compression, which tackles spatio-temporal redundancy in dynamic sequences; and (3) audio-centric compression, which handles temporal and spectral redundancy in acoustic signals. Beyond this modality-driven categorization, we further dissect methods based on their underlying mechanisms, including transformation-based, similarity-based, attention-based, and query-based approaches. By providing a comprehensive and structured overview, this survey aims to consolidate current progress, identify key challenges, and inspire future research directions in this rapidly evolving domain. For ongoing updates and to track the latest advances in this promising area, we maintain a public repository: https://github.com/cokeshao/Awesome-Multimodal-Token-Compression.
@@
Multimodal Large Language Models, Token Compression, Long Context, Efficient AI1 Introduction
Multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] have demonstrated exceptional performance on complex tasks, including visual question answering (VQA), automatic speech recognition (ASR) and multimodal content generation, by extending the architectural principles of large language models (LLMs) [12, 13, 14, 15]. These powerful models derive their strength from processing very long and diverse contexts, such as high-resolution images, extended video sequences, and long audio input with the transformer architectures.
Achieving this capability, however, face a significant challenge: the quadratic complexity of the self-attention mechanism. As the number of tokens increases, this complexity leads to substantial computational and memory demands. This problem is particularly pronounced in MLLMs, where the tokenization of visual and audio data can generate sequences of orders of magnitude longer than text [16, 17, 18].
For instance, as illustrated in Figure 1, the number of image tokens is directly proportional to resolution, while the number of video tokens scales with both resolution and duration, and audio tokens are proportional to duration. A single content-rich video can produce tens of millions of tokens, dramatically exacerbating computational inefficiencies and leading to severe inference latency (90 minutes video will be converted into 54M tokens)111. Consequently, addressing this computational bottleneck is critical for unlocking the full potential of MLLMs in real-world applications.
@@
Audio-centric token compression mitigates temporal and spectral redundancy, as salient information often concentrates within sparse, brief segments and specific frequency bands amidst long silent pauses or stationary background noise.
@@
The remaining sections of the article are organized as follows: we will first discuss the architecture of MLLMs in the background section (Section 2.1), followed by an examination of how token compression has been utilized in prior methods for large language models (LLMs, Section 2.2) and vision transformers (ViTs, Section 2.3) Subsequent sections will be dedicated to token compression methods for specific modalities: Section 3 for image LLMs, Section 4 for video LLMs, and Section 5 for audio LLMs. Following this, Section 6 will provide insights into token compression research. Finally, Section 7 will introduce the broad application space of token compression, followed by the concluding Section 8.
@@
2 Background
2.1 Multimodal Architecture
@@
The high dimensionality of multimodal data and the resulting computational demands remain a challenge. As shown in Figure 2, the token sequence processed comprises a mix of system prompt, multimodal context, textual instruction. In most reasoning tasks, multimodal tokens can account for over of the total sequence length, thereby forming the primary computational bottleneck. This bottleneck is a critical obstacle to scaling MLLMs and achieving efficient inference. Consequently, a key strategy for optimizing computational efficiency involves employing specialized projector architectures. These projectors are designed to reduce the number of multimodal tokens while preserving their semantic fidelity, thus mitigating the computational burden.
While the MLLM architecture presents unique challenges, token compression research has been explored for both encoders and LLMs independently. Therefore, the subsequent sections will first dive into token compression techniques relevant to these individual components, paving the way for more efficient multimodal models. Specifically, Section 2.2 will focus on token compression methods for large language models (LLMs), and Section 2.3 will explore techniques for vision transformers (ViTs).
2.2 Large Language Model Token Compression
@@
AutoCompressor [133] condenses context into summary vectors as soft prompts. Extensible Tokenization [134] employs intermediate modules to compress embeddings, while SentenceVAE [135] represents sentences with single tokens. Selective Context [136] employs self-information metrics to eliminate low-information tokens. LLMLingua [137, 138, 139] series utilizes hierarchical token pruning with instruction tuning and further introduces LongLLMLingua [138] to mitigate position decay through semantic density ranking. In parallel, query-guided methods like QUITO [140] and QUITO-X [141] leverage attention scores or information bottleneck theory for relevance-based filtering. AdaComp [142] implements adaptive extraction governed by query complexity predictors. Concept Distillation [143] employs Abstract Meaning Representation (AMR) graphs to distill key concepts, whereas xRAG [144] collapses documents into single-token representations. ICAE [145] encodes context into discrete memory slots. Recursive frameworks including RCC [146] and XL3M [147] generate piecewise summaries through relevant fusion. Additionally, SoftPromptComp [148] fuses natural language prompts with dynamic embeddings, while PromptIntern [149] internalizes task instructions into model parameters via phased training.
@@
2.3 Vision Transformer Token Compression
@@
| Cons | Limited by the transformation method, compression rate isn’t flexible enough | May lose fine-grained info if tokens over-generalized; poor structural feature retention if token differentiation is low | Explicit attention score calculation might be incompatible with mainstream acceleration libraries | Not user-friendly for multi-turn conversations; requires re-condensing information |
3 Image-centric Token Compression
Multimodal long context token compression methods generally fall into four categories based on their underlying mechanisms: transformation-based approaches directly transform the cross-modality information to compress tokens by altering their scale or representation; similarity-based techniques reduce tokens by leveraging the inherent resemblances between them; attention-based strategies exploit the sparsity of attention within the multimodal data to guide compression; and query-based methods selectively refine multimodal information, guided by prompts, to distill the most irrelevant tokens. Each of these methods has its own set of advantages and disadvantages, which are summarized in Table I. Representative image-centric token compression methods are further compared in Table II.
3.1 Transformation-based Image-centric Compression
@@ 3.1.1 Pixel Unshuffle
3.1.2 Spatial Pooling / Interpolation
@@ 3.1.3 Spatial Convolution
3.1.4 Comparative Analysis of Transformation Methods
These transformation-based image-centric compression methods effectively utilize all image tokens while consciously preserving the spatial local information of 2D features. Pixel unshuffle, pooling, and interpolation are inherently parameter-free, thus introducing no additional weight overhead, which is a key advantage. In contrast, methods that use convolution introduce trainable weights to learn a more sophisticated local abstraction.
@@ 3.1.4 Comparative Analysis of Transformation Methods
3.2 Similarity-based Image-centric Compression
@@
Early works in this area include ToMe [60], an acceleration method for ViTs. ToMe introduces a token merge module between the attention and MLP blocks, calculating token similarity and merging similar tokens via bipartite soft matching. In the context of MLLMs, FOLDER [61] employs a similar approach, inserting a token merge module within the last attention block of the vision encoder. This reduces the number of tokens that were subsequently passed to the LLM decoder. DivPrune [28] reframes the token compression problem as a Max-Min diversity problem [161], aiming to select a subset of tokens with maximal internal differences. AuroraCap [45] adopts a strategy consistent with ToMe, performing token merging within each attention and MLP block of the vision tower. This progressively reduces the number of tokens throughout the ViT model. While the aforementioned methods primarily leverage similarity-based clustering of tokens within the ViT, TopV [49] extends this principle to compress tokens within the LLM layers. TopV comprehensively considers both the similarity and distance functions between features to guide the token compression process, operating directly within the multimodal representation space of the LLM.
3.3 Attention-based Image-centric Compression
@@ 3.3.2 Attention in Decoder
3.3.3 Critical Challenge for Pruning in Decoder
While these methods leverage attention scores within the LLM decoder to offer sophisticated ways to compress visual tokens, they face a significant practical challenge: the explicit need to access attention scores. This direct access is often incompatible with highly optimized acceleration libraries like FlashAttention [163, 164], which compute attention implicitly or in a fused manner for speed. This incompatibility can be mitigated by performing an additional, separate attention calculation solely for pruning purposes. However, for multi-layer pruning strategies such as FitPrune, ST3, and ZipVL, this additional computational overhead becomes significantly more pronounced, potentially negating the efficiency gains.
3.4 Query-based Image-centric Compression
@@
3.4.1 Token Distillation
Token distillation originates from the early projector designs of MLLMs. The goal is to distill visual tokens to learn the most text-relevant visual representations, thereby reducing visual tokens while simultaneously aligning modalities.
@@ 3.4.1 Token Distillation
3.4.2 Cross-Modal Selection
@@
4 Video-centric Token Compression
@@
To address this, current video LLMs commonly employ a 1 FPS sampling rate to reduce token counts. Furthermore, unlike the typical approach for single images, where the image is split into patches and fed into an encoder for detailed information, video processing often foregoes this detailed frame-level segmentation to keep token numbers manageable. Even with these strategies, the quantity of video tokens remains substantial. During model training and understanding, transformation-based methods, such as the pooling technique used in LLaVA-Video [34], are frequently employed to reduce tokens and aid the model’s comprehension of video content.
@@
To fully grasp token compression for video LLMs, it is highly recommended to first review Section 3, which details spatial compression methods. Next, we will primarily discuss techniques addressing the temporal domain. Similar to image-centric methods, selected video-centric token compression methods are compared in Table III.
4.1 Transformation-based Video-centric Compression
Like image LLMs, video LLMs use image encoders for visual tokens. Consequently, transformation-based video-centric compression methods fundamentally operate on the principles established in Section 3.1, with the added capability of performing 3D transformations. A multitude of models showcase cross-modal applicability, performing effectively in both image and video inference tasks. Following the structure of Section 3.1, we will now detail transformation-based video-centric compression methods.
4.1.1 2D/3D Pooling
@@ 4.1 Transformation-based Video-centric Compression
This dual focus on spatial and increasingly on temporal pooling underscores their combined importance in enabling efficient processing and comprehensive understanding of video content, particularly as video durations extend. SlowFast-LLaVA-1.5 [92] incorporates a two-stream SlowFast projector into a LLaVA-style architecture, using a slow pathway to sample fewer, spatially rich frames and a fast pathway to sample more, spatially compressed frames, then concatenates both for the LLM—achieving efficient long-form video understanding with reduced token count while preserving spatiotemporal details.
4.1.2 2D/3D Convolution
Similar to pooling, convolution can also be employed for downsampling video tokens, but it does so in a parameterized manner. Instead of simply aggregating information like pooling, convolution layers learn filters to process and condense spatial and temporal features. VideoLLaMA 2 [10], for instance, thoroughly investigated both 2D and 3D pooling and convolution approaches. Their experiments ultimately showed that 3D convolution yielded the best balance of performance and efficiency for video token downsampling. This suggests that learning intricate spatiotemporal relationships through convolutions is more effective for comprehensive video understanding compared to pooling alone.
4.2 Similarity-based Video-centric Compression
Given the temporal redundancy inherent in video, where adjacent frames often exhibit high visual similarity, temporal compression is frequently prioritized over, or integrated with spatial compression. To effectively handle this temporal redundancy, video frames are typically first clustered.
@@ 4.2 Similarity-based Video-centric Compression
While some methods do not explicitly cluster video frames, FrameFusion [79], for example, acts as a token compression technique for streaming video LLMs. It directly merges temporally redundant tokens exceeding a specific threshold in the shallow layers of the model.
4.3 Attention-based Video-centric Compression
Current attention-based token compression methods in video LLMs and image LLMs share significant similarities. When attention is applied within the encoder to guide token compression, videos are typically treated as a sequence of independent images fed into an image encoder. Consequently, attention computations largely disregard inter-frame relationships, making these approaches fundamentally image-centric token compression. For a more concise discussion of such attention-based methods, please refer to Section 3.3.
In contrast, methods employing attention within the decoder process video frames sequentially, concatenating their tokens over time. For longer videos, particularly in the context of streaming video LLMs, windowed attention is commonly used to mitigate computational overhead by focusing on local temporal visual information. However, it’s notable that even these windowed attention-based methods within the decoder often share the same foundational principles as those discussed in Section 3.3.
4.4 Query-based Video-centric Compression
4.4.1 Token Distillation

### v4

@@ Abstract
Multimodal large language models (MLLMs) have made remarkable strides, largely driven by their ability to process increasingly long and complex contexts, such as high-resolution images, extended video sequences, and lengthy audio input. While this ability significantly enhances MLLM capabilities, it introduces substantial computational challenges, primarily due to the quadratic complexity of self-attention mechanisms with numerous input tokens. To mitigate these bottlenecks, token compression has emerged as an auspicious and critical approach, efficiently reducing the number of tokens during both training and inference. In this paper, we present the first systematic survey and synthesis of the burgeoning field of multimodal long context token compression. Recognizing that effective compression strategies are deeply tied to the unique characteristics and redundancies of each modality, we categorize existing approaches by their primary data focus, enabling researchers to quickly access and learn methods tailored to their specific area of interest: (1) image-centric compression, which addresses spatial redundancy in visual data; (2) video-centric compression, which tackles spatio-temporal redundancy in dynamic sequences; and (3) audio-centric compression, which handles temporal and spectral redundancy in acoustic signals. Beyond this modality-driven categorization, we further dissect methods based on their underlying mechanisms, including transformation-based, similarity-based, attention-based, and query-based approaches. By providing a comprehensive and structured overview, this survey aims to consolidate current progress, identify key challenges, and inspire future research directions in this rapidly evolving domain.
@@
Multimodal Large Language Models, Token Compression, Long Context, Efficient AII Introduction
Multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] have demonstrated exceptional performance on complex tasks, including visual question answering (VQA), automatic speech recognition (ASR) and multimodal content generation, by extending the architectural principles of large language models (LLMs) [12, 13, 14, 15]. These powerful models derive their strength from processing very long and diverse contexts, such as high-resolution images, extended video sequences, and long audio input, using transformer architectures.
Achieving this capability, however, faces a significant challenge: the quadratic complexity of the self-attention mechanism. As the number of tokens increases, this complexity leads to substantial computational and memory demands. This problem is particularly pronounced in MLLMs, where the tokenization of visual and audio data can generate sequences of orders of magnitude longer than text [16, 17, 18].
For instance, as illustrated in Figure 1, the number of image tokens is directly proportional to resolution, while the number of audio tokens is proportional to duration, and video tokens scale with both resolution and duration. A single content-rich video can produce tens of millions of tokens, dramatically exacerbating computational inefficiencies and leading to severe inference latency (90 minutes video will be converted into 54M tokens)111. Consequently, addressing this computational bottleneck is critical for unlocking the full potential of MLLMs in real-world applications.
@@
Audio-centric token compression mitigates temporal and spectral redundancy, as salient information often concentrates within sparse, brief segments and specific frequency bands amidst silent pauses or background noise.
@@
The remaining sections of the article are organized as follows: we will first discuss the architecture of MLLMs in the background section (Section II-A), followed by an examination of how token compression has been utilized in prior methods for large language models (LLMs, Section II-B) and vision transformers (ViTs, Section II-C). Subsequent sections will be dedicated to token compression methods for specific modalities: Section III for image LLMs, Section IV for video LLMs, and Section V for audio LLMs. Following this, Section VI will provide insights into token compression research. Finally, Section VII will introduce the broad application space of token compression, followed by the concluding Section VIII.
@@
II Background
II-A Multimodal Architecture
@@
The high dimensionality of multimodal data poses a computational challenge. As shown in Figure 2, the token sequence processed comprises a mix of system prompt, multimodal context, and textual instruction. In most reasoning tasks, multimodal tokens constitute over of the total sequence length [26], thereby forming the primary computational bottleneck. This bottleneck obstacle to scaling MLLMs and achieving efficient inference. Consequently, a key strategy to optimize computational efficiency involves employing specialized projector architectures. These projectors are designed to reduce the number of multimodal tokens while preserving their semantic fidelity, thus mitigating the computational burden.
While MLLM architecture presents unique challenges, token compression has been explored for both encoders and LLMs independently. Therefore, the subsequent sections will first dive into techniques relevant to these individual components, paving the way for more efficient multimodal models. Specifically, Section II-B will focus on token compression methods for large language models (LLMs), and Section II-C will explore techniques for vision transformers (ViTs).
II-B Large Language Model Token Compression
@@
AutoCompressor [133] condenses context into summary vectors as soft prompts. Extensible Tokenization [134] employs intermediate modules to compress embeddings, while SentenceVAE [135] represents sentences with single tokens. Selective Context [136] employs self-information metrics to eliminate low-information tokens. LLMLingua [137, 138, 139] series utilizes hierarchical token pruning with instruction tuning and further introduces LongLLMLingua [138] to mitigate position decay through semantic density ranking. In parallel, query-guided methods like QUITO [140] and QUITO-X [141] leverage attention scores or information bottleneck theory for relevance-based filtering. AdaComp [142] implements adaptive extraction governed by query complexity predictors. Concept Distillation [143] employs Abstract Meaning Representation (AMR) graphs to distill key concepts, whereas xRAG [144] collapses documents into single-token representations. ICAE [145] encodes context into discrete memory slots. Recursive frameworks including RCC [146] and XL3M [147] generate piecewise summaries through relevant fusion. SoftPromptComp [148] fuses natural language prompts with dynamic embeddings, while PromptIntern [149] internalizes task instructions into model parameters via phased training.
@@
II-C Vision Transformer Token Compression
@@
| Cons | Limited by the transformation method, compression rate isn’t flexible enough | May lose fine-grained info if tokens over-generalized; poor structural feature retention | Explicit attention score calculation might be incompatible with mainstream acceleration libraries | Not user-friendly for multi-turn conversations; requires re-condensing information |
III Image-centric Token Compression
Multimodal long context token compression methods generally fall into four categories based on their underlying mechanisms: transformation-based approaches directly transform the cross-modality information to compress tokens by altering their scale or representation; similarity-based techniques reduce tokens by leveraging the inherent resemblances between them; attention-based strategies exploit the sparsity of attention within the multimodal data to guide compression; and query-based methods selectively refine multimodal information, guided by prompts, to distill the most relevant tokens. Each of these methods has its own set of advantages and disadvantages, which are summarized in Table I. Representative image-centric token compression methods are further compared in Table II.
III-A Transformation-based Image-centric Compression
@@ III-A1 Pixel Unshuffle
III-A2 Spatial Pooling / Interpolation
@@ III-A3 Spatial Convolution
III-A4 Comparative Analysis of Transformation Methods
These transformation-based image-centric compression methods effectively utilize all image tokens while consciously preserving the spatial local information of 2D features. Pixel unshuffle, pooling, and interpolation are inherently parameter-free, thus introducing no additional weight overhead, a key advantage. In contrast, convolution learn a more sophisticated local abstraction by introducing trainable weights.
@@ III-A4 Comparative Analysis of Transformation Methods
III-B Similarity-based Image-centric Compression
@@
Early works in this area include ToMe [60], an acceleration method for ViTs. ToMe introduces a token merge module between the attention and MLP blocks, calculating token similarity and merging similar tokens via bipartite soft matching. This process creates a new set of tokens, , by replacing the most similar tokens with their merged representations.
where each is a set of tokens identified as highly similar by ToMe’s matching algorithm.
In the context of MLLMs, FOLDER [61] employs a similar approach, inserting a token merge module within the last attention block of the vision encoder. This reduces the number of tokens that were subsequently passed to the LLM decoder. DivPrune [28] reframes the token compression problem as a Max-Min diversity problem [161], aiming to select a subset of tokens with maximal internal differences. AuroraCap [45] adopts a strategy consistent with ToMe, performing token merging within each attention and MLP block of the vision tower. This progressively reduces the number of tokens throughout the ViT model. While the aforementioned methods primarily leverage similarity-based clustering of tokens within the ViT, TopV [49] extends this principle to compress tokens within the LLM layers. TopV comprehensively considers both the similarity and distance functions between features to guide the token compression process, operating directly within the multimodal representation space of the LLM.
III-C Attention-based Image-centric Compression
@@ III-C2 Attention in Decoder
III-C3 Critical Challenge for Pruning in Decoder
While these methods leverage attention scores within the LLM decoder to offer sophisticated ways to compress visual tokens, they face a significant practical challenge: the explicit need to access attention scores. This direct access is often incompatible with highly optimized acceleration libraries like FlashAttention [163, 164], which compute attention implicitly or in a fused manner for speed. This incompatibility can be mitigated by performing an additional, separate attention calculation solely for pruning purposes. However, for progressive pruning strategies such as FitPrune, ST3, and ZipVL, this additional computational overhead becomes significantly more pronounced, potentially negating the efficiency gains.
III-D Query-based Image-centric Compression
@@
III-D1 Token Distillation
Token distillation originates from the early projector designs of MLLM. The goal is to distill visual tokens to learn the most text-relevant visual representations, reduce visual tokens while simultaneously align modalities.
@@ III-D1 Token Distillation
III-D2 Cross-Modal Selection
@@
IV Video-centric Token Compression
@@
To address this, current video LLMs commonly employ a 1 FPS sampling rate to reduce token counts. Furthermore, unlike methods for single images, which often encode both the global image and a series of local patches for detailed feature extraction, video processing often foregoes this detailed frame-level segmentation to keep token numbers manageable. Even with these strategies, the quantity of video tokens remains substantial. During model training and understanding, transformation-based methods, such as the pooling technique used in LLaVA-Video [34], are usually employed to reduce tokens and aid the model’s comprehension of video content.
@@
To fully grasp token compression for video LLMs, it is highly recommended to first review Section III, which details spatial compression methods. Next, we will primarily discuss techniques addressing the temporal domain. Similar to image-centric methods, selected video-centric token compression methods are compared in Table III.
IV-A Transformation-based Video-centric Compression
Like image LLMs, video LLMs use encoders for visual tokens. Consequently, transformation-based video-centric compression methods fundamentally operate on the principles established in Section III-A, with the added capability of performing 3D transformations. A multitude of models showcase cross-modal applicability, performing effectively in both image and video inference tasks. Following the structure of Section III-A, we will now detail transformation-based video-centric compression methods.
IV-A1 2D/3D Pooling
@@ IV-A Transformation-based Video-centric Compression
This dual focus on spatial and increasingly on temporal pooling underscores their combined importance in enabling efficient processing and comprehensive understanding of video content, particularly as video durations extend. SlowFast-LLaVA [92] incorporates a two-stream SlowFast projector into a LLaVA-style architecture, using a slow pathway to sample fewer, spatially rich frames and a fast pathway to sample more, spatially compressed frames, then concatenates both for the LLM—achieving efficient long-form video understanding with reduced token count while preserving spatiotemporal details.
IV-A2 2D/3D Convolution
Similar to pooling, convolution can also be employed for downsampling video tokens, but it does so in a parameterized manner. Instead of simply aggregating information like pooling, convolution layers learn filters to process and condense spatial and temporal features. VideoLLaMA 2 [10], for instance, thoroughly investigated both 2D and 3D pooling and convolution approaches. Their experiments showed that 3D convolution yielded the best balance of performance and efficiency for video token downsampling. This suggests that learning intricate spatiotemporal relationships through convolutions is more effective for comprehensive video understanding compared to pooling alone.
IV-B Similarity-based Video-centric Compression
Given the temporal redundancy inherent in video, where adjacent frames often exhibit high visual similarity, temporal compression is frequently prioritized over or integrated with spatial compression. To effectively handle this temporal redundancy, video frames are typically first clustered.
@@ IV-B Similarity-based Video-centric Compression
While some methods do not explicitly cluster video frames, FrameFusion [79], for example, acts as a token compression technique for video LLMs. It directly merges temporally redundant tokens exceeding a specific threshold in the shallow layers of the model.
IV-C Attention-based Video-centric Compression
Current attention-based token compression methods in video LLMs and image LLMs share significant similarities. When attention is applied within the encoder to guide token compression, videos are typically treated as a sequence of images fed into an image encoder, making these approaches similar to image-centric token compression. For a more concise discussion of such attention-based methods, please refer to Section III-C.
In contrast, methods employing attention within the decoder process video frames sequentially, concatenating their tokens over time. For longer videos, particularly in the context of streaming video LLMs, windowed attention is commonly used to mitigate computational overhead by focusing on local temporal visual information. However, it’s notable that even these windowed attention-based methods within the decoder often share the same foundational principles as those discussed in Section III-C.
IV-D Query-based Video-centric Compression
IV-D1 Token Distillation